OpenStudy (anonymous):

find 2 additional points that lie on the line that passes through the point (-4,3) and slope m=1/3

OpenStudy (anonymous):

to do this we have to know the equation. to find the equation we use the point slope form that is y-y1 = m(x-x1). I know that looks confusing but that y1 and x1 is just from the coordinates you provided. so putting x1 and y1 in their places(don't forget that if any of the points are negative that we have to compensate for that when we replace them (y - -1 = y+1) so that is y - 3 = (1/3)(x+4) we solve for y = max + b because that is the simplest way to look at a line and that gives us y = (1/3)x +13/3.

OpenStudy (anonymous):

Ok so now that we have that we can substitute different points in the equation. lets just make it simple by finding the x-intercept, where y = 0. so we have 0 = (1/3)x + 13/3. (13/3 = 4.33333 if you want to do it that way, which i will from now on). we isolate the x by minusing 4.33333 from both sides and get -4.3333 = (1/3)x and multiply both sides by 3/1(which is just 3) to cancel out the 1/3. so x = -4.33333 * 3 = 13. (either by calculator or by (13/3)*(3/1) the two threes cancel out and just leave 13). and then substitue something like 1 in for x. and you have your two points. :)

OpenStudy (anonymous):

my mistake, that one that i worked out was (13,0) and x = 1 would be (1/3) + (13/3) = 14/3. so (1, 14/3)